OpenStudy (nali):

use the power-reducing formulas to rewrite the expression in terms of the first power of the cosine cos^4 2x

OpenStudy (nali):

iam confused about the first part so can someone just show me hoe to do the beginning and then i will continue

OpenStudy (anonymous):

\[ \cos^2(x)= \frac 1 2 ( 1+ \cos(2x)) \]

OpenStudy (anonymous):

\[ \cos^4(x)= \frac 1 4 (1+\cos(2x))^2 \]

OpenStudy (nali):

ok after that what do i do to change ^^ into co^4 2x

OpenStudy (anonymous):

\[ (\cos (2 x)+1)^2=\cos ^2(2 x)+2 \cos (2 x)+1 \]

OpenStudy (anonymous):

\[ \cos^2(2 x) = \frac 1 2( 1 + \cos(4x)) \]

OpenStudy (anonymous):

Putting them together, you get \[\cos ^4(x)=\frac{1}{8} (4 \cos (2 x)+\cos (4 x)+3) \]

OpenStudy (anonymous):

Did you get it?

OpenStudy (nali):

but iam looking for cos^4 2x

OpenStudy (nali):

not cos^4 x

OpenStudy (anonymous):

Replace every x by 2x in the answer and you will get \[ \cos ^4(2 x)=\frac{1}{8} (4 \cos (4 x)+\cos (8 x)+3) \]

OpenStudy (nali):

thanku so much

OpenStudy (anonymous):

yw

OpenStudy (nali):

by the way if i close this question and i want to check it again can it or will it be deleted by tomorrow?

OpenStudy (anonymous):

No it will not be deleted.

OpenStudy (nali):

so i could check at any time in the closed questions section

OpenStudy (anonymous):

I think so.

OpenStudy (nali):

ok thanks again